OpenStudy (carissa15):

I have a few Chemistry questions, I have tried to go through them but keep getting stuck. Can anyone help me please?

OpenStudy (carissa15):

The first two I have are precipitate questions. I am not sure whether the distilled water "topping" up the volume makes any difference to my calculations? for the first one I Keep getting 7.599g BaCrO4 which is the precipitate? But perhaps I am not factoring the amount that is used (if any in the reaction). So lost....

OpenStudy (frostbite):

Where do you get the 0.06 from? Solved that from the equilibrium condition?

OpenStudy (carissa15):

\[0.03 mol K2CrO4 (\frac{ 2 mol KCl }{ 1 mol K2CrO4 }) =0.06\]

OpenStudy (aaronq):

He's referring to the concentration for KCl.

OpenStudy (carissa15):

I used that 0.06 mol KCl and since the volume is 1L had concentration for KCl as 0.06M. Is this right?

OpenStudy (aaronq):

if you had 100% conversion from potassium chromate, but you're not using the limiting reactant which is barium chloride.

OpenStudy (aaronq):

So the most possible moles of barium chromate made was 0.02 moles. Now how much remains dissolved and how much is solid? The Ksp equation gives you the solubility. So 1.45*10^(-5) M, the moles of barium chromate dissolved are \(\sf M=\dfrac{moles}{L}\rightarrow moles=molarity*L=1.45*10^{-5} M*1L=1.45*10^{-5} moles\) Now, we subtract whats dissolved from the total: 0.02-1.45*10^(-5)=0.0199855086232538 moles Multiply by the molar mass of barium chromate, the mass is 5.0637 g

OpenStudy (carissa15):

Thankyou. So just to clarify. As BaCl2 was the limiting that is why the BaCrO4 has to be 0.02 moles also? Then, the rest of the calculations I had done work. Thank you so much :)

OpenStudy (carissa15):

I am stuck on another precipitation question but this time I need to calculate the concentration of the magnesium ions remaining in solution?

OpenStudy (aaronq):

The 0.02 moles of BaCl2 and 0.02 moles of barium chromate would not exist simultaneously. BaCl2 breaks apart in solution and the barium ion finds chromate. The most possible (theoretical) amount of barium chromate you could make is related to the moles of the limiting reactant, you have to factor out the coefficients. Have you ever done an I.C.E. table? Not that you need it, but it illustrates the concept.

OpenStudy (carissa15):

I have done ICE tables but get mixed up with what is the change and what is initial etc

OpenStudy (aaronq):

It's a similar concept. Write the Ksp and use the \(\sf [OH^-] \) obtained from the pH.

OpenStudy (aaronq):

(that was for the second question)

OpenStudy (aaronq):

Hm you should work with the ICE table concept, there are tons of examples on youtube, or we could do one here if you'd like - i think the visuals (youtube) are more effecitve.

OpenStudy (carissa15):

cool, I will watch some and get my head around the concept :)

OpenStudy (aaronq):

awesome! do ask again if you're still confused

OpenStudy (carissa15):

Could you please help walk me through the PH Question?

OpenStudy (carissa15):

So I put Ksp = [Na] [OH] and then use Na as X and OH as PH 12.0?

OpenStudy (aaronq):

sure. now that i'm doing it, i dont think you need the pH at all - since it's only asking for the concentration remaining in solution - which is given by the Ksp. The trick here is writing the Ksp equation, we need the balanced equation for the process, can you provide that?

OpenStudy (aaronq):

ps. if you want to find \(\sf [OH^-]\) from pH, you need to find pOH first. pH + pOH= 14 pOH=14-pH \(\sf pOH=-log[OH^-]\)

OpenStudy (carissa15):

I had NaFl + CaCl = CaFl +NaCl

OpenStudy (carissa15):

Handy to remember :)

OpenStudy (aaronq):

:)

OpenStudy (aaronq):

um thats not the right equation :S

OpenStudy (carissa15):

oh oops..

OpenStudy (aaronq):

\(\sf NaOH + MgX_2 \rightleftharpoons Mg(OH)_2 +NaX\) X is just a soluble counter-ion

OpenStudy (aaronq):

can you balance it?

OpenStudy (carissa15):

2Na + MgX2 = Mg(OH)2 +2NaX

OpenStudy (carissa15):

2NaOH***

OpenStudy (aaronq):

okay cool. Now we need the net ionic equation, can you do that?

OpenStudy (carissa15):

Is it 2OH +Mg =Mg(OH)2 as the Na and X are the spectators?

OpenStudy (aaronq):

yup, you need to show the charges of the ions. the Ksp equation is: \(\sf K_{sp}=(2[OH^-])^2*[Mg^{2+}]\) we let \(\sf [OH^-]=[Mg]^{2+}=x\) then we have, \(\sf Ksp=(2x^2)*x=4x^3\)

OpenStudy (aaronq):

I have to get going, you should repost this question if you still need help with it!

OpenStudy (carissa15):

Thank you so much for your help :)

OpenStudy (aaronq):

no problem :)